A volume of 450cm³ of air is measured at a pressure of 740mm Hg absolute and a temperature of 20°C. What is the volume in cm³ at 760 mm Hg absolute and 0°C?


1
Expert's answer
2021-06-07T15:01:35-0400

The ideal gas equation of state says


"\frac{p_1V_1}{T_1}=\frac{p_2V_2}{T_2}"

Hence, the final volume


"V_2=\frac{p_1V_1T_2}{p_2T_1}=\frac{740*450*273}{760*293}=408\:\rm cm^3"